Suppose these are the market demand and supply curves for hooded sweatshirts:
Supply: P = 10 + 2QS Demand: P = 50−3QD

(a) Sketch these two curves (that is, draw them, but don’t worry about numerical accuracy). Calculate equilibrium price and quantity. Calculate equilibrium price and quantity.

(b) Show on your graph the areas of consumer and producer surplus. Calculate consumer and producer surplus at the equilibrium from part a.

(c) Calculate the price elasticity of demand when price changes from $29 to $26, and from $26 to $23. Is demand elastic or inelastic over those two regions?

(d) Calculate elasticity of supply as price changes from $26 to $28. Is supply elastic or inelastic?
Answer #1

6 price 07 t - - - - - o 2 4 6 8 10 12 14 16 18 20 Quantity p=10+2 05 P 10 20 p=50-32 [ppsation are In Equilibrium D=5so-3a = 10+20 So 10 – 20+30 4o - sa Q = no S a = 8 unit Substituting a into supply equation p = 10+2 (8) P =10+16 p =26 (6) c(C) P.S = 64 p=50-3Qd on P = २१ २१ = 50-321 300 = 50-२१ २५ २ -१ when p=26 २८ -50-321 34 =5०-२६ २] = २५ 24 = 8 units. ११ - २५Ed - & 8-7 8t7 26-29 26729 - 3 1 27.5 7.5 27.5 - - - / s x-27.5 3 = -27.5 22.5 Ed = -1.22 demand is since Ed = 1.221. Hence E

when p=28 28 = 10+2 25 205 = 28 to Qs = 18 2 Qs = q units. «Ese P26 P2=28 Es = 9 8 28-26 9t8 28+26 2 2 8.S 27 17 Es = 1.58 >